Joe has an equal number of nickles and dimes. If he has $1.05 in all, how many coins of each type does he have?
Found 2 solutions by josgarithmetic, addingup:
Answer by josgarithmetic(39799)
Answer by addingup(3677)

n = d
0.05n+0.10d = 1.05 substitute for n
0.05(d)+0.10d = 1.05
0.05d+0.10d = 1.05
0.15d = 1.05
d = 7 he has 7 dimes, and since he has an equal number of nickels, he has 7 nickels
-----------------------------
Check:
7*0.10 = 0.70
7*0.05 = 0.35
Total . . 1.05 Correct
:
John